# Realised Volatility Forecasting: Machine Learning via Financial Word Embedding

### v5

@@ A.1 Data Cleaning Details
The text cleaning procedures are grouped into five main categories: 1) Primary, 2) Begins with, 3) Ends with, 4) General, and 5) Final checks. The Primary rules focus on fundamental transformations of the raw data, including extracting the body of news from extensible markup language (XML), removing XML-encoding characters (XMLENCOD), converting XML content into plain text through parsing, transforming upper-case letters into lower-case, and removing embedded tables. These steps ensure that the core textual content is isolated and normalised before applying more specific pattern-based cleaning rules. This category forms the foundation of the cleaning pipeline and prepares the text for further refinement.
@@ A.5 HAR-Family of Models: In-Sample and Out-of-Sample Results
AR1 HAR HAR-J CHAR SHAR ARQ HARQ HARQ-F $\beta_{0}$ 3.6436a 1.8528 1.4401 1.2273 1.8444 2.8734 1.6127 0.6807 3.6436b 1.8530 1.4408 1.2479 1.8447 2.8734 1.6135 0.8452 (1.6305)c (0.9216) (0.6446) (0.6357) (0.9213) (1.3585) (0.8758) (0.8382) (100, 0)d (100, 0) (99.9, 0.1) (98.7, 1.3) (99.9, 0.1) (100, 0) (99.9, 0.1) (77.2, 22.8) $\beta_{1}$ 0.1428 0.0394 0.6544 0.3653 0.2136 0.1695 0.1428 0.0421 0.6544 0.3653 0.2136 0.1702 (0.1090) (0.0382) (0.2743) (0.2028) (0.1093) (0.1193) (100, 0) (93.8, 6.2) (100, 0) (100, 0) (100, 0) (99.3, 0.7) $\beta^{+}_{1}$ 0.0625 0.0638 (0.0942) (92.8, 7.2) $\beta^{-}_{1}$ 0.0346 0.0397 (0.0390) (88.4, 11.6) $\beta^{w}$ 0.1471 0.0523 0.1449 0.1123 0.2076 0.1642 0.0925 0.1616 0.1341 0.2197 (0.1805) (0.1305) (0.1774) (0.1568) (0.1731) (77.2, 22.8) (60.2, 39.8) (77.0, 23.0) (72.6, 27.4) (89.6, 10.4) $\beta^{m}$ 0.3566 0.2510 0.3527 0.3090 0.5367 0.3566 0.2529 0.3530 0.3097 0.5368 (0.1754) (0.1562) (0.1767) (0.1675) (0.1956) (100, 0) (97.1, 2.9) (98.7, 1.3) (98.1, 1.9) (100, 0) $\beta_{jump}$ -0.7353 0.7353 (0.2840) (0, 100) $\beta_{BPV}^{d}$ 0.4136 0.4161 (0.2526) (98.3, 1.7) $\beta^{w}_{BPV}$ 0.3434 0.3697 (0.2926) (91.2, 8.8) $\beta^{m}_{BPV}$ 0.4336 0.4433 (0.3329) (92.5, 7.5) $\beta_{Q}$ -0.0004 -0.0003 -0.0002 0.0004 0.0003 0.0002 (0.0004) (0.0003) (0.0003) (0.4, 99.6) (0.6, 99.4) (2.0, 98.0) $\beta^{w}_{Q}$ -0.0008 0.0011 (0.0013) (22.0, 78.0) $\beta^{m}_{Q}$ -0.0064 0.0068 (0.0080) (7.4, 92.6) $Adj.\ R^{2}(avg)$ 0.0318 0.0721 0.0984 0.1000 0.0731 0.0583 0.0818 0.0871 $Adj.\ R^{2}(med)$ 0.0104 0.0404 0.0673 0.0692 0.0406 0.0406 0.0505 0.0552 MSE(avg)MSE\ (avg) 173.3761 165.6939 161.0308 160.7043 165.4823 168.4723 163.9610 162.7886 MSE(med)MSE\ (med) 127.6872 124.0916 120.5695 119.9260 124.0427 125.2517 121.4500 120.8497 QLIKE(avg)QLIKE\ (avg) 0.6139 0.5619 0.5177 0.5167 0.5615 0.5773 0.5481 0.5599 QLIKE(med)QLIKE\ (med) 0.6223 0.5670 0.5187 0.5236 0.5656 0.5813 0.5452 0.5737 Notes: For each parameter in each model, the $\text{mean of coefficients}^{[a]}$, the $\text{mean of absolute coefficients}^{[b]}$, the $\text{standard deviation}^{[c]}$, and the proportions of positive and negative $\text{coefficients}^{[d]}$ are computed using 36,892 daily estimations (1,604 days $\times$ 23 stocks). ‘HAR-J’ is HAR with jump, ‘CHAR’ stands for continuous HAR, and ‘SHAR’ denotes semivariance HAR. ‘ARQ’, ‘HARQ’, and ‘HARQ-F’ are the introduced models in Bollerslev et al. (2016). $\beta_{0}$ is the constant term, $\beta_{1}$ is the coefficient of the first lag of RV, $\beta^{+}_{1}$ and $\beta^{-}_{1}$ are the coefficients of the first lag of RV for the positive and negative returns, $\beta^{w}$ and $\beta^{m}$ are the coefficients of the daily average of RV over the last week and last month, $\beta_{jump}$ is the coefficient of the jump term of ‘HAR-J’ model, $\beta_{BPV}^{d}$, $\beta^{w}_{BPV}$, and $\beta^{m}_{BPV}$ are the coefficients of continuous terms of the ‘CHAR’ model, and $\beta_{Q}$, $\beta^{w}_{Q}$, and $\beta^{m}_{Q}$ stand for the coefficients of the first lag of RQ, and the daily average of RQ for the last week and last month, respectively. In the lower section of this table, bolded values represent the highest performance for each metric. For each stock, $MSE$ and $QLIKE$ are initially calculated as the average (or median) in-sample errors. These values are subsequently aggregated across all 23 stocks, using either the overall average (or median).

### v6

@@ A.1 Data Cleaning Details
The text cleaning procedures are grouped into five main categories: 1) Primary, 2) Begins with, 3) Ends with, 4) General, and 5) Final checks. The Primary rules focus on fundamental transformations of the raw data, including extracting the body of news from extensible markup language (XML), removing XML-encoding characters (XMLENCOD), converting XML content into plain text through parsing, converting upper-case letters into lower-case, and removing embedded tables. These steps ensure that the core textual content is isolated and normalised before applying more specific pattern-based cleaning rules. This category forms the foundation of the cleaning pipeline and prepares the text for further refinement.
@@ A.5 HAR-Family of Models: In-Sample and Out-of-Sample Results
AR1 HAR HAR-J CHAR SHAR ARQ HARQ HARQ-F $\beta_{0}$ 3.6436a 1.8528 1.4401 1.2273 1.8444 2.8734 1.6127 0.6807 3.6436b 1.8530 1.4408 1.2479 1.8447 2.8734 1.6135 0.8452 (1.6305)c (0.9216) (0.6446) (0.6357) (0.9213) (1.3585) (0.8758) (0.8382) (100, 0)d (100, 0) (99.9, 0.1) (98.7, 1.3) (99.9, 0.1) (100, 0) (99.9, 0.1) (77.2, 22.8) $\beta_{1}$ 0.1428 0.0394 0.6544 0.3653 0.2136 0.1695 0.1428 0.0421 0.6544 0.3653 0.2136 0.1702 (0.1090) (0.0382) (0.2743) (0.2028) (0.1093) (0.1193) (100, 0) (93.8, 6.2) (100, 0) (100, 0) (100, 0) (99.3, 0.7) $\beta^{+}_{1}$ 0.0625 0.0638 (0.0942) (92.8, 7.2) $\beta^{-}_{1}$ 0.0346 0.0397 (0.0390) (88.4, 11.6) $\beta^{w}$ 0.1471 0.0523 0.1449 0.1123 0.2076 0.1642 0.0925 0.1616 0.1341 0.2197 (0.1805) (0.1305) (0.1774) (0.1568) (0.1731) (77.2, 22.8) (60.2, 39.8) (77.0, 23.0) (72.6, 27.4) (89.6, 10.4) $\beta^{m}$ 0.3566 0.2510 0.3527 0.3090 0.5367 0.3566 0.2529 0.3530 0.3097 0.5368 (0.1754) (0.1562) (0.1767) (0.1675) (0.1956) (100, 0) (97.1, 2.9) (98.7, 1.3) (98.1, 1.9) (100, 0) $\beta_{jump}$ -0.7353 0.7353 (0.2840) (0, 100) $\beta_{BPV}^{d}$ 0.4136 0.4161 (0.2526) (98.3, 1.7) $\beta^{w}_{BPV}$ 0.3434 0.3697 (0.2926) (91.2, 8.8) $\beta^{m}_{BPV}$ 0.4336 0.4433 (0.3329) (92.5, 7.5) $\beta_{Q}$ -0.0004 -0.0003 -0.0002 0.0004 0.0003 0.0002 (0.0004) (0.0003) (0.0003) (0.4, 99.6) (0.6, 99.4) (2.0, 98.0) $\beta^{w}_{Q}$ -0.0008 0.0011 (0.0013) (22.0, 78.0) $\beta^{m}_{Q}$ -0.0064 0.0068 (0.0080) (7.4, 92.6) $Adj.\ R^{2}(avg)$ 0.0318 0.0721 0.0984 0.1000 0.0731 0.0583 0.0818 0.0871 $Adj.\ R^{2}(med)$ 0.0104 0.0404 0.0673 0.0692 0.0406 0.0406 0.0505 0.0552 MSE(avg)MSE\ (avg) 173.3761 165.6939 161.0308 160.7043 165.4823 168.4723 163.9610 162.7886 MSE(med)MSE\ (med) 127.6872 124.0916 120.5695 119.9260 124.0427 125.2517 121.4500 120.8497 QLIKE(avg)QLIKE\ (avg) 0.6139 0.5619 0.5177 0.5167 0.5615 0.5773 0.5481 0.5599 QLIKE(med)QLIKE\ (med) 0.6223 0.5670 0.5187 0.5236 0.5656 0.5813 0.5452 0.5737 Notes: For each parameter in each model, the $\text{mean of coefficients}^{[a]}$, the $\text{mean of absolute coefficients}^{[b]}$, the $\text{standard deviation}^{[c]}$, and the proportions of positive and negative $\text{coefficients}^{[d]}$ are computed using 36,892 daily estimations (1,604 days $\times$ 23 stocks). ‘HAR-J’ is HAR with jump, ‘CHAR’ stands for continuous HAR, and ‘SHAR’ denotes semivariance HAR. ‘ARQ’, ‘HARQ’, and ‘HARQ-F’ are the introduced models in Bollerslev et al. (2016). $\beta_{0}$ is the constant term, $\beta_{1}$ is the coefficient of the first lag of RV, $\beta^{+}_{1}$ and $\beta^{-}_{1}$ are the coefficients of the first lag of RV for the positive and negative returns, $\beta^{w}$ and $\beta^{m}$ are the coefficients of the daily average of RV over the last week and last month, $\beta_{jump}$ is the coefficient of the jump term of ‘HAR-J’ model, $\beta_{BPV}^{d}$, $\beta^{w}_{BPV}$, and $\beta^{m}_{BPV}$ are the coefficients of continuous terms of the ‘CHAR’ model, and $\beta_{Q}$, $\beta^{w}_{Q}$, and $\beta^{m}_{Q}$ stand for the coefficients of the first lag of RQ, and the daily average of RQ for the last week and last month, respectively. For each stock, $MSE$ and $QLIKE$ are initially calculated as the average (or median) in-sample errors. These values are subsequently aggregated across all 23 stocks, using either the overall average (or median).